# Self-supervised Learning for Geospatial AI: A Survey

### v1

@@ VI-A Multi-type Learning
While SSL techniques have demonstrated promising performance in GeoAI, these methods focus on separately deriving representations for individual geospatial data types. This paradigm does not consider the complex interactions and potential synergies among various geospatial data types. To this end, recent studies propose multi-type learning methods, which involve the joint learning of representations for multiple geospatial data types, as a step towards the development of geospatial foundation models. For example, several studies [166, 167, 168] approach this goal by utilizing a variety of entities such as POIs, regions, and their associated content like street view images, remote sensing images, and user data, to construct an urban knowledge graph. The edges between these entities are established by considering factors such as topological relations, human mobility patterns, semantic relations, and similarities in POI distributions across regions. Then knowledge graph embedding techniques are applied to derive the representations for multiple geospatial data types.
@@
VI-B Adaptation to Language Space
The integration of LLMs, which are fundamentally rooted in the SSL paradigm, represents a burgeoning trend across various domains, including GeoAI. LLMs have demonstrated not only enhanced language understanding, reasoning, and generation capabilities, but also remarkable adaptability to various other data modalities, including images [52, 171], graphs [172, 173], and time series [174, 175]. Furthermore, LLMs have been validated to possess a certain level of knowledge in geospatial domain [176, 177, 178]. Based on these capabilities, several studies have started to explore the application of LLMs to geospatial data types within language space for GeoAI. These efforts typically employ alignment strategies to facilitate language-based interactions with geospatial data through targeted SSL fine-tuning of LLMs.
@@ VI-B Adaptation to Language Space
GeoLLM [179] proposes harness the capabilities of LLMs to encode geospatial knowledge and capture regional features effectively. Specifically, it pinpoints targeted region coordinates on the map, extracts the corresponding address that contains place names from the neighborhood level up to the country, and identifies the ten nearest places. The regional coordinates with its geospatial context are then formatted into templated textual prompts as input to the LLM. After processing the prompts, the LLM is designed to be automatically fine-tuned using response variables, such as various socio-economic indicators. LAMP [180] aims to infuse fine-grained knowledge about POI into LLMs for a specific city, subsequently facilitating POI-related applications in a conversational manner. To achieve this, this method structures several POI search tasks with their ground truth data as templates within the SSL corpus to fine-tune the LLM. Consequently, LAMP can solve several POI applications, such as route recommendation and location search, by formulating them as query-response processes within LLM framework. UrbanGPT [181] introduces a SSL instruction-tuning paradigm that seeks to align the dependencies of time and space, with the language space of LLMs. This method constructs prompts that combine textual descriptions with representations obtained from spatio-temporal learning models for LLM. The resulting output representations encapsulate both semantic information and relevant time-space dependencies for geospatial regions.
VI-C Task-specific SSL techniques
Although the primary focus of this survey is on deriving generalizable representations of various geospatial data types for broad application across multiple downstream tasks, there are numerous studies that employ more task-specific SSL techniques for pre-training or as supplementary objectives to enhance the main task. These specialized approaches tailor SSL techniques to meet the unique demands and intricacies of specific geospatial applications.
@@ VI-C Task-specific SSL techniques
A prominent example of task-specific SSL is in spatio-temporal forecasting tasks. For example, models like UniST [182], W-MAE [183], and GPT-ST [184] are utilized for grid regions or sensor points, employing a strategy of reconstructing masked features as self-supervised pre-training method to learn dynamic dependencies. The learned parameters are then fine-tuned for specific spatio-temporal forecasting datasets. Besides, SSL techniques are also integrated within multi-task learning frameworks for spatio-temporal forecasting. UrbanSTC [185] applies contrastive learning to identify grid regions in both spatial and temporal dimension with similar patterns. STGCL [186] explores various contrastive learning schemes with in the framework of spatio-temporal GNN, such as node-level and graph-level contrasts, providing some insights into effective integration strategies. ST-SSL [187] performs the adaptive augmentation over the traffic flow graph at both attribute- and structure-levels within the spatio-temporal GNN framework. It introduces two SSL auxiliary tasks to supplement the main traffic forecasting task to account for spatial and temporal heterogeneity. SSTBAN [188] incorporates a masked auto-encoder module to reconstruct masked spatio-temporal patches, thus deriving more robust representations to support the forecasting task. CL4ST [189] develops a meta view generator to automatically construct node and edge augmentation views for contrastive learning in a data-driven manner. Moreover, several studies also incorporate contrastive SSL into the scenario of POI recommendation [190, 191, 192]. These diverse methods demonstrate the versatility and potential of task-specific SSL techniques in enhancing the performance for various downstream applications regarding geospatial data.
VII Future Research Directions
@@
Selection of pretext tasks and data augmentation. Pretext tasks and data augmentation techniques play a crucial role in the effectiveness of SSL. Existing SSL methods in GeoAI usually draw inspiration from the domains of CV and graph learning, employing heuristic adaptations tailored to geospatial contexts. As a result, the selection of pretext tasks and data augmentation strategies can vary significantly across different geospatial SSL implementations. While there have been efforts to systematically assess the effectiveness of combining different pretext tasks and data augmentation techniques in other domains [193, 194, 195], these results may not directly translate to geospatial SSL due to the unique characteristics and diversity of geospatial data types. Therefore, there is a pressing need to rigorously investigate the impact of different pretext task and data augmentation selections specifically within the geospatial domain. This investigation should ideally be supported by theoretical analysis and supplemented by comprehensive empirical evaluations to ensure robustness. Moreover, it is beneficial to construct standard benchmarks for geospatial SSL. This would not only facilitate the comparison of different SSL approaches but also provide valuable guidance for the development of future research efforts in various downstream applications.
@@
Multi-modality data fusion. With the rapid expansion of location-based services and spatial crowdsourcing, data from diverse sources attached to geospatial objects, previously difficult to access, becomes increasingly available, including street view images, textual comments, and videos. While there are several efforts to integrate multi-modality data to enhance the performance of geospatial applications [196, 197, 198], only few have explored the ideas of SSL [92, 91]. Integrating multi-modality data sources for geospatial objects presents both challenges and opportunities. Models designed for multi-modality data fusion must handle the discrepancies in scale, resolution, and relevance across different data sources. Based on this, future research can explore novel architectures, pretext tasks, and fusion techniques that effectively leverage the complementary information from different data sources. For example, adapting models like CLIP [52] with geospatial awareness. It would also be interesting to study how to balance the contributions of different modalities, aiming to achieve more generalizable and robust representations in real-world scenarios.
@@
Geospatial foundation models and LLM adaptation. As discussed in Section VI, LLMs and multi-type pre-trained models opens up exciting possibilities for adaptation to the geospatial context, serving as a universal basis for various geospatial applications. In terms of pathways in pre-training geospatial foundation models, future research could focus on creating SSL techniques that can be efficiently trained on massive-scale geospatial datasets [199, 200]. These models could learn generalized representations of geospatial features, patterns, and relationships, forming a foundation to be fine-tuned for specific tasks. On the other hand, adapting LLMs to the geospatial context requires novel learning paradigms that align spatial relationships and geographic context within the language space. For instance, there is a critical need to develop cross-modal training techniques that effectively bridge textual, visual, and spatial data.
@@
Privacy and vulnerability in geospatial SSL. The inherent nature of geospatial data, which often contains sensitive information about individuals and urban infrastructure, raises significant privacy and data vulnerability concerns. For instance, aggregated trajectories, when analyzed with certain approaches, can expose individual’s routes and private locations [201]. To this end, privacy-preserving techniques, such as differential privacy, could be considered within the SSL framework. Moreover, federated learning approaches [202, 203] can be explored, which enable the collaborative training of models on sensitive geospatial data while avoiding the direct sharing of raw data, thus maintaining privacy. Furthermore, geospatial SSL models, like other machine learning models, are susceptible to adversarial attacks where input data is manipulated to induce model errors, which are particularly problematic in urban decision-making process. Therefore, robust SSL techniques can be developed to resist such data poisoning attacks or adversarial examples [204, 205].
Dynamic model learning and updating. Most existing SSL techniques for geospatial data employ the setting of learning static representations. However, as urban environments continuously evolve, the characteristics of geospatial objects may experience distribution shifts. Therefore, it is desirable to develop more flexible SSL techniques that possess the ability of dynamically updating in response to new environmental conditions. One promising direction is the integration of continual learning mechanisms into SSL techniques [206]. This would allow models to incorporate new information and adapt to changes without the need for complete retraining, while also preventing catastrophic forgetting of previously learned knowledge. Furthermore, investigating the active learning principles with SSL techniques could lead to more efficient updating strategies [207]. By identifying and utilizing the most informative new samples for model adaptation, these methods could significantly reduce the computational and data demands for model updating.
VIII Conclusion
@@
This paper provides a comprehensive overview of self-supervised learning in the domain of GeoAI. We develop a structured framework and introduce a systematic taxonomy that organizes SSL based on three geospatial data types and two methodology categories. For each data type, we offer detailed descriptions of methods, summarize their key features within the SSL component, and discuss their downstream applications. We further present studies on the emerging trends for GeoAI and task-specific SSL techniques. Finally, we outline several promising directions for the research in the future. As this domain continues to expand and evolve, we hope that the discussion presented in this paper will contribute to the future advancements in GeoAI.

### v2

@@ 6 Multi-type Learning
While SSL techniques have demonstrated promising performance in producing representations for geospatial objects, these methods focus on separately deriving representations for individual geospatial data types. This paradigm does not consider the complex interactions and potential synergies among various geospatial data types. To this end, recent studies propose multi-type learning methods, which involve the joint modeling for multiple geospatial data types, as a step towards the development of geospatial foundation models.
One research direction involves the construction of geospatial knowledge graphs to systematically represent various geospatial objects as nodes and their relationships with diverse data sources as edges. By structuring geospatial entities within a graph-based framework, knowledge graphs enable the encoding of complex spatial, semantic, and contextual relationships among various data types. Once constructed, graph embedding techniques can be applied to derive representations that capture both the intrinsic properties of geospatial objects and their interactions within the broader spatial environment. WorldKG[185] aligns geospatial objects from OpenStreetMap with ontologies in existing knowledge graphs such as Wikidata. It structures geospatial knowledge by defining graph nodes based on geospatial objects, semantic classes, and object attributes. In contrast, frameworks such as UUKG [186], UrbanKG [187], and UrbanFlood [188] develop customized graph schemas tailored to specific research perspectives, such as user aspect [187], hierarchical spatial resolution [186], or application-oriented [188] geospatial analysis. Following the construction of these knowledge graphs, various graph embedding techniques [189] can be employed to generate representations for each node. These embeddings facilitate downstream geospatial tasks by leveraging the rich structural and semantic information encoded within the knowledge graph framework from multiple data types.
@@
7 Discussion
Sections 3-6 have presented specialized SSL techniques designed for different geospatial data types. Through an extensive review of prior studies, we observe a notable evolution in SSL methodologies for geospatial representation learning, characterized by the following key trends:
  1. 1.
    Transition from Single-Source to Multi-Source Context Integration: Early research predominantly derived representations from a single source of context information. However, recent advancements increasingly emphasize multi-source and multi-modal integration, leveraging diverse and complementary context signals to better capture intrinsic spatial patterns in geospatial data.
  2. 2.
    Transition from Naive Models to Hybrid Architectures: Initial approaches typically employ simple models (e.g., Word-2vec or RNN variants) for geospatial representation learning. Over time, research has shifted towards hybrid models that combine multiple and complex architectural components, enabling more expressive modeling of geospatial context information.
  3. 3.
    Influence of Foundation Models from Visual and Language Domains: The rapid progress in visual and language foundation models has significantly influenced geospatial representation learning. Recent studies increasingly incorporate these foundation models as integral components or explore methods to align the implementation of these pre-trained models in the context of geospatial data.
While the studies belonging to the first two research trends have been well-established and detailed in preceding sections, the emergence of geospatial foundation models remains an evolving area with no fixed formulation and is explored from various perspectives. Recognizing this new emerging trend, we provide an overview of recent efforts in developing geospatial foundation models from multiple perspectives.
Additionally, while the primary focus of this survey is to summarize SSL-based studies that adopt a task-agnostic approach to geospatial representation learning, another important research direction involves task-specific SSL techniques. Unlike general-purpose SSL, these methods employ self-supervised objectives as auxiliary learning signals to enhance supervised tasks in geospatial applications. To offer a broader perspective, we also present a brief overview of research efforts that incorporate task-specific SSL into several typical geospatial applications.
7.1 A Step Towards Geospatial Foundation Models
Foundation models, which are fundamentally rooted in the SSL paradigm with pretext tasks such as autoregressive modeling in language domains and masked patch recovery in vision domains, have emerged as a transformative trend across various domains, including GeoAI. Foundation models have demonstrated significant advancements in understanding, reasoning, and generative capabilities across diverse data modalities, including language [43], images [51, 192], graphs [193, 194], and time series [195, 196]. Inspired by the success, researchers have increasingly explored the development of geospatial foundation models primarily based on language models from various perspectives.
One perspective involves enhancing existing foundation models with geospatial knowledge. Empirical studies have shown that LLMs, such as the GPT series, already possess a certain level of geospatial knowledge [197, 198, 199]. To further strengthen their geospatial knowledge understanding, researchers have proposed the development of geospatial language models [200, 201, 202] by fine-tuning general-purpose LLMs within an autoregressive SSL paradigm. These models are trained on domain-specific corpora, including research publications and Wikipedia pages related to geography, urban planning, and spatial sciences, enabling them to perform better on geospatial tasks such as geospatial question answering.
Beyond continued fine-tuning, another perspective explores leveraging LLMs for complex geospatial problem-solving through planning, decision-making, and pipeline orchestration. LLMs are increasingly utilized as intelligent agents capable of organizing systematic workflows for geospatial tasks. This is achieved through task decomposition, strategic tool selection, and model integration, where LLMs interact with external spatiotemporal models and analytical tools to process queries and generate informed responses. Notable examples include UrbanLLM [203], GeoGPT [204], PlanGPT [205], and MapGPT [206], which are designed to answer geospatial queries by organizing the workflow of solving these queries by calling external functions (e.g., ArcGIS) and integrating outputs from them. While these approaches contribute to the broader development of geospatial foundation models, they are often not explicitly designed for direct representation learning of geospatial objects.
@@ 7.1 A Step Towards Geospatial Foundation Models
Moreover, several studies have explored to enhance the interaction between LLMs and particular geospatial objects through targeted alignment strategies by SSL. For example, LAMP [207] aims to infuse fine-grained knowledge about POI into LLMs for a specific city, subsequently facilitating POI-related applications in a conversational manner. To achieve this, it structures several POI search tasks with their ground truth data as templates within the SSL corpus to fine-tune the LLM. Consequently, LAMP can solve several POI applications, such as route recommendation and location search, by formulating them as query-response processes within LLM framework. For polyline object, TrajFM [208] develops a trajectory modeling framework that pre-trains LLMs from scratch in an autoregressive manner, integrating spatial, temporal, and POI modalities for each trajectory data point. To enhance task transferability, TrajFM employs a trajectory masking and recovery scheme, which unifies various trajectory-related task generation processes by masking and reconstructing trajectory sub-segments and modalities. This enables the pre-trained model to generalize across diverse trajectory-based tasks. TrajCogn [209] transforms trajectory features into structured, natural language-like inputs by mapping these features into descriptive words, allowing LLMs to process trajectory data effectively. The model is then fine-tuned using a cross-reconstruction pretext task, enabling generalizable utility for downstream applications such as travel time estimation, destination prediction, and trajectory similarity search. For polygon object, GeoLLM [210] proposes harness the capabilities of LLMs to encode geospatial knowledge and capture regional features effectively. Specifically, it pinpoints targeted region coordinates on the map, extracts the corresponding address that contains place names from the neighborhood level up to the country, and identifies the ten nearest places. The regional coordinates with its geospatial context are then formatted into templated textual prompts as input to the LLM. After processing the prompts, the LLM is designed to be automatically fine-tuned using response variables, such as various socio-economic indicators. UrbanGPT [211] introduces a SSL instruction-tuning paradigm that seeks to align the dependencies of time and space, with the language space of LLMs. This method constructs prompts that combine textual descriptions with representations obtained from spatio-temporal learning models for LLM. The resulting output representations encapsulate both semantic information and relevant time-space dependencies for geospatial regions.
7.2 Task-specific SSL techniques
As previously discussed, task-specific SSL techniques serve as auxiliary objectives to enhance performance in specific geospatial tasks. These methods are particularly relevant for data instances such as grids (rasters) and point sensors, where various types of measurements (e.g., flows, speeds) are recorded over time, forming a collection of time series from these instances that exhibit both spatial and temporal interdependencies.
@@ 7.2 Task-specific SSL techniques
Therefore, a notable application of task-specific SSL applied to these data instances is in the application of spatio-temporal forecasting based on the collected records. For example, models like UniST [155], W-MAE [156], and GPT-ST [212] are utilized for grids or sensor points, employing a strategy of reconstructing masked features as self-supervised pre-training method to learn dynamic dependencies. The learned parameters are then fine-tuned for specific spatio-temporal forecasting datasets. Besides, SSL techniques are also integrated within multi-task learning frameworks for spatio-temporal forecasting. UrbanSTC [213] applies contrastive learning to identify grid regions in both spatial and temporal dimension with similar patterns. STGCL [214] explores various contrastive learning schemes with in the framework of spatio-temporal GNN, such as node-level and graph-level contrasts, providing some insights into effective integration strategies. ST-SSL [215] performs the adaptive augmentation over the traffic flow graph at both attribute- and structure-levels within the spatio-temporal GNN framework. It introduces two SSL auxiliary tasks to supplement the main traffic forecasting task to account for spatial and temporal heterogeneity. SSTBAN [216] incorporates a masked auto-encoder module to reconstruct masked spatio-temporal patches, thus deriving more robust representations to support the forecasting task. CL4ST [217] develops a meta view generator to automatically construct node and edge augmentation views for contrastive learning in a data-driven manner. Moreover, several studies also incorporate contrastive SSL into the scenario of POI recommendation [218, 219, 220]. These diverse methods demonstrate the versatility and potential of task-specific SSL techniques in enhancing the performance for various downstream applications regarding geospatial data.
8 Future Research Directions
@@
Selection of pretext tasks and data augmentation. Pretext tasks and data augmentation techniques play a crucial role in the effectiveness of SSL. Existing SSL methods for geospatial objects usually draw inspiration from the domains of computer vision and graph learning, employing heuristic adaptations tailored to geospatial contexts. As a result, the selection of pretext tasks and data augmentation strategies can vary significantly across different geospatial SSL implementations. While there have been efforts to systematically assess the effectiveness of different pretext tasks and data augmentation techniques in other domains [221, 222, 223], these findings may not directly translate to geospatial SSL due to the unique characteristics and diversity of geospatial data types. Therefore, there is a pressing need to investigate the impact of different pretext task and data augmentation selections specifically within the geospatial domain. Future research should explore whether certain pretext tasks or augmentation techniques offer distinct advantages for geospatial representation learning. This inquiry should ideally be guided by theoretical analyses and comprehensive empirical evaluations to ensure the robustness and generalizability of SSL techniques for geospatial objects.
Benchmarking SSL for geospatial objects. The establishment of standardized benchmarks is essential for advancing SSL techniques in geospatial domain. Unlike domains such as computer vision or natural language processing, where benchmark datasets are widely available and standardized, SSL methods for geospatial objects are currently evaluated on datasets that are often task-specific and unique to individual studies. The absence of unified benchmarking datasets hinders consistent and reliable comparisons across different methods, making it difficult to systematically assess their effectiveness and generalizability. Future research should prioritize the development of well-structured benchmarking frameworks that facilitate comprehensive evaluations of SSL models for geospatial objects. This requires a concerted effort to curate and release open-access benchmark datasets [224] that encompass diverse geospatial data types, heterogeneous contextual information, and multiple downstream tasks. Additionally, the establishment of automated benchmarking platforms would significantly advance this domain by enabling standardized evaluations of SSL models. Such platforms could provide pre-configured training pipe-lines, fair comparison protocols, and leaderboards, similar to benchmark suites in other AI disciplines.
@@
Enhanced multi-modality fusion. With the rapid expansion of location-based services and spatial crowdsourcing, data from diverse sources attached to geospatial objects, previously difficult to access, becomes increasingly available, including street view images, textual comments, and videos. Several efforts have been made to integrate multi-modality models to enhance the performance of geospatial applications [225, 226, 227]. However, there remains significant untapped potential in advancing model capabilities for geospatial tasks. Integrating multi-modality data sources for geospatial objects presents both challenges and opportunities. Models designed for multi-modality data fusion must handle the discrepancies in scale, resolution, and relevance across different data sources. Future research can explore novel architectures, pretext tasks, and fusion techniques that effectively leverage the complementary information from different data sources. For example, adapting models like CLIP [51] with geospatial awareness. It would also be interesting to study how to balance the contributions of different modalities, aiming to achieve more generalizable and robust representations in real-world scenarios.
@@
Geospatial foundation models and LLM adaptation. As discussed in Section 7, LLMs and multi-type pre-trained models opens up exciting possibilities for adaptation to the geospatial context, serving as a universal basis for various geospatial applications. In terms of pathways in pre-training geospatial foundation models, future research could focus on creating SSL techniques that can be efficiently trained on massive-scale geospatial datasets [228, 229]. These models could learn generalized representations of geospatial features, patterns, and relationships, forming a foundation to be fine-tuned for specific tasks. On the other hand, adapting LLMs to the geospatial context requires novel learning paradigms that align spatial relationships and geographic context within the language space. For instance, there is a critical need to develop cross-modal training techniques that effectively bridge textual, visual, and spatial data for enhancing the applicability in more diverse scenarios and tasks in multiple modalities.
@@
Privacy and vulnerability. The inherent nature of geospatial data, which often contains sensitive information about individuals and urban infrastructure, raises significant privacy and data vulnerability concerns [230]. For instance, aggregated trajectories, when analyzed with certain approaches, can expose individual’s routes and private locations [231]. To this end, privacy-preserving techniques, such as differential privacy, could be considered within the SSL framework, especially the pre-training corpus. Moreover, federated learning approaches [232, 233] can be explored, which enable the collaborative training of models on sensitive geospatial data while avoiding the direct sharing of raw data, thus maintaining privacy. Furthermore, geospatial SSL models, like other machine learning models, are susceptible to adversarial attacks where input data is manipulated to induce model errors, which are particularly problematic in urban decision-making process. Therefore, robust SSL techniques can be developed to resist such data poisoning attacks or adversarial examples [234, 235].
9 Conclusion
@@
This paper provides a comprehensive overview of self-super-vised learning for geospatial objects in the domain of GeoAI. We develop a structured framework and introduce a systematic taxonomy that organizes SSL based on three geospatial data types and two methodology categories. For each data type, we offer detailed descriptions of methods, summarize their key features within the SSL component, and discuss their downstream applications, along with the utilization of multi-type SSL techniques. We further present studies on the emerging trends and task-specific SSL techniques. Finally, we outline several promising directions for the research in the future. As this domain continues to expand and evolve, we hope that the discussion presented in this paper will contribute to the future advancements of GeoAI.